# Stable Offline Value Function Learning with Bisimulation-based Representations

### v2

@@ C.2 Baselines
We provide details of the baselines in this section.
@@
The dr3 regularizer minimizes the total feature co-adaptation by adding the term $\sum_{(s,a,s^{\prime})\in\mathcal{D},a^{\prime}\sim{\pi_{e}}}\phi(s,a)^{\top}\phi(s^{\prime},a^{\prime})$ as an auxiliary task to the main fqe loss (Kumar et al., 2021). Ma et al. (2024) introduced an improvement to this auxiliary loss by suggesting that the absolute value of the feature co-adaptation be minimized, i.e., $\sum_{(s,a,s^{\prime})\in\mathcal{D},a^{\prime}\sim{\pi_{e}}}|\phi(s,a)^{\top}\phi(s^{\prime},a^{\prime})|$. e use $\alpha=0.1$ as its auxiliary task weight. Absolute dr3’s only hyperparameters are the auxiliary task weight $\alpha$ and the $\phi$ output dimension.
@@
He et al. (2024) introduced an alternative regularizer to dr3 rank regularizer since they suggested that the minimization of the unbounded feature co-adaptation can undermine performance. They introduced their bounded rank regularizer beer (see Equation (12) in He et al. (2024)). beer introduces only the auxiliary task weight $\alpha$ as the additional hyperparameter.
@@ C.3.1 Understanding the Stability of the krope Learning Procedure
We conduct our experiments on the Markov reward process in Figure 5(a), which was introduced by Feng et al. (2019). The mrp consists of $4$ non-terminal states, $1$ terminal state (the box), and only $1$ action. The value function estimate is linear in the weight $w=[w_{1},w_{2},w_{3}]$, so the native features of the states are $[1,0,0]$, $[0,1,0]$ , $[0,0,2]$, and $[0,0,1]$ (clockwise starting from left-most). We set $\gamma=1$, so the optimal weights are $w^{*}=[0.8,1,0]$. In this setup, we say a transition is a bad transition if the bootstrapping target is a moving target for the current state (Asadi et al., 2024). For example, the transition from $w_{3}$ to $2w_{3}$ is a bad transition since updates made to $w_{3}$ may move $2w_{3}$ further away. When this transition is sampled at a frequency that is different from the on-policy distribution, algorithms such as td, lspe, and fqe tend to diverge (Asadi et al., 2024).
@@
C.3.2 Offline Policy Evaluation on d4rl Datasets
In this section, we present the offline policy evaluation results on the d4rl datasets. The setup is the same as the one used in Section 4. We present the results in Figure 6.
@@ C.3.2 Offline Policy Evaluation on d4rl Datasets
Qualitatively, we reach the similar conclusions: krope is effective in producing stable and accurate ope estimates. However, in $3/9$ instances, krope does diverge. This divergence is likely related to the discussion in Section 5 and Appendix C.3.1. Recall that krope is a semi-gradient method, which does not optimize any objective function and is susceptible to divergence (Feng et al., 2019; Sutton & Barto, 2018). So while krope representations stabilize value function learning, krope’s learning algorithm may diverge and not converge to krope representations. However, regardless of this result, krope does improve the stability and accuracy of fqe in all cases.
@@ C.3.3 Stability-related Analysis on Custom Datasets
Feature Co-adaptation, Condition Number, and Hyperparameter Sensitivity.
@@
(c) WalkerStand
Figure 7: Feature co-adaptation on different environments as a function of training epochs. All results are averaged over $20$ trials and shaded region is the $95\%$ confidence interval.
(a) CheetahRun
(b) FingerEasy
(c) WalkerStand
Figure 8: Condition number of the covariance matrix on different environments as a function of training epochs. All results are averaged over $20$ trials and shaded region is the $95\%$ confidence interval.
(a) CheetahRun
(b) FingerEasy
(c) WalkerStand
@@ C.3.4 Using fqe Directly for ope
For the sake of completeness, we present results of fqe as an ope algorithm where the action-value network is directly used to estimate the performance of ${\pi_{e}}$. We present the results in Figures 10 and  11. As done in Section 4, we evaluate the performance of fqe and krope based on how they shape the penultimate features of the action-value network. However, when conducting ope, we evaluate two variants: 1) using lspe (-l) and 2) using the same end-to-end fqe action-value network (-e2e).
@@
Regardless of fqe’s hyperparameter sensitivity, it is still interesting to observe that when fqe is used as an ope algorithm, it produces reasonably accurate ope estimates. It even outperforms the fqe+krope combination. However, as shown in Section 4, the penultimate features of this same network actually have poor property values such as high feature co-adaptation, high condition numbers, and highly erratic ope estimates with lspe. Given that these features have weak properties, it is unclear why fqe still can lead to accurate ope. The primary difference between using fqe for ope vs. fqe features and lspe for ope is in how the last linear layer is trained. The former is trained by gradient descent while the latter is trained with the iterative lspe algorithm on the fixed features. An interesting future direction will be to explore the learning dynamics of these two approaches.
On a related note, we point out that the training dynamics of fqe are still not well-understood. For example, Fujimoto et al. (2022) show that the fqe loss function poorly correlates with value error. That is, the fqe loss can be high but value error (and ope error) can be low.

### v3

@@ C.2 Baselines
We provide details of the $6$ baselines in this section.
dbc.
It is a bisimulation-based metric learning algorithm that learns $\pi$-bisimilar representations by ensuring the L1111 distance between the latent representations approximates the $\pi$-bisimulation metric (Zhang et al., 2021). During the learning process, it also learns a model of the environment. Its additional hyperparameters are the learning rate of model learning (which we set to $1e^{-5}$) and the output dimension of $\phi$.
rope.
It is a bisimulation-based metric learning algorithm that is off-policy evaluation variant of mico (Castro et al., 2022; Pavse & Hanna, 2023a). It learns representations directly such that the $\pi$-bisimulation metric is satisfied and makes no assumptions on the transition dynamics. Its additional learning rate is the output dimension of $\phi$.
@@
The dr3 regularizer minimizes the total feature co-adaptation by adding the term $\sum_{(s,a,s^{\prime})\in\mathcal{D},a^{\prime}\sim{\pi_{e}}}\phi(s,a)^{\top}\phi(s^{\prime},a^{\prime})$ as an auxiliary task to the main fqe loss (Kumar et al., 2021). (Ma et al., 2024) introduced an improvement to this auxiliary loss by suggesting that the absolute value of the feature co-adaptation be minimized, i.e., $\sum_{(s,a,s^{\prime})\in\mathcal{D},a^{\prime}\sim{\pi_{e}}}|\phi(s,a)^{\top}\phi(s^{\prime},a^{\prime})|$. We use $\alpha=0.1$ as its auxiliary task weight. Absolute dr3’s only hyperparameters are the auxiliary task weight $\alpha$ and the $\phi$ output dimension.
@@
(He et al., 2024) introduced an alternative regularizer to dr3 rank regularizer since they suggested that the minimization of the unbounded feature co-adaptation can undermine performance. They introduced their bounded rank regularizer beer (see Equation (12) in (He et al., 2024)). beer introduces only the auxiliary task weight $\alpha$ as the additional hyperparameter.
@@ C.3 Additional Results
(a) Realizability Error
Figure 4: Evaluation of basic representation properties on Garnet mdps with $40$ state-actions vs. output dimension $d$. Figure 4(a): Realizability error; lower is better. All results are averaged over $30$ trials and the shaded region is the $95\%$ confidence interval.
Realizability.
In addition to stability, we also care about the realizability of $\Phi$. We say $\Phi$ is a realizable representation if $q^{{\pi_{e}}}\in\text{Span}(\Phi),$ where $\text{Span}(\Phi)$ is the subspace of all expressible action-value functions with $\Phi$. Note that even if $\Phi$ is a realizable and stable representation, lspe may not recover the $q^{{\pi_{e}}}$ solution (Sutton & Barto, 2018).
A basic criterion for learning $q^{\pi_{e}}$ is realizability. That is, we want ${\epsilon}:=\|\Phi\hat{w}-q^{{\pi_{e}}}\|^{2}_{2}$, where $\hat{w}:=\operatorname*{arg\,min}_{w}\|\Phi w-q^{{\pi_{e}}}\|_{2}^{2}$, to be low. In our experiments, we compute ${\epsilon}$ and plot it as a function of $d$ in Figure 4(a). A critical message from our results is that stability and realizability do not always go hand-in-hand. While bcrl-exp-na has favorable spectral radius properties (Figure 2(a)), it has poor realizability, which will negatively affect its ope accuracy. krope, on the other hand, has favorable stability and realizability properties up till $d=30$. When $d\geq 40$, the realizability error is 00 for all algorithms since the subspace spanned by $\Phi$ is large enough to contain the true action-value function (Ghosh & Bellemare, 2020). While the realizability error is 00 for $d\geq 40$, the representations can be unstable (Figure 2(a)).
@@ C.3.1 Understanding the Stability of the krope Learning Procedure
We conduct our experiments on the Markov reward process in Figure 5(a), which was introduced by (Feng et al., 2019). The mrp consists of $4$ non-terminal states, $1$ terminal state (the box), and only $1$ action. The value function estimate is linear in the weight $w=[w_{1},w_{2},w_{3}]$, so the native features of the states are $[1,0,0]$, $[0,1,0]$ , $[0,0,2]$, and $[0,0,1]$ (clockwise starting from left-most). We set $\gamma=1$, so the optimal weights are $w^{*}=[0.8,1,0]$. In this setup, we say a transition is a bad transition if the bootstrapping target is a moving target for the current state (Asadi et al., 2024). For example, the transition from $w_{3}$ to $2w_{3}$ is a bad transition since updates made to $w_{3}$ may move $2w_{3}$ further away. When this transition is sampled at a frequency that is different from the on-policy distribution, algorithms such as td, lspe, and fqe tend to diverge (Asadi et al., 2024).
@@
C.3.2 Learning Graphs for Offline Policy Evaluation on dmc and d4rl Datasets
In this section, we present the offline policy evaluation results on the dmc and d4rl datasets. We present the results in Figures 6 and 7.
(a) CartPoleSwingUp
(b) CheetahRun
(c) FingerEasy
(d) WalkerStand
Figure 6: Normalized squared value error achieved by lspe when using a particular representation vs. representation training epochs on the custom dmc datasets. lspe estimates are computed every $10$ epochs. Results are averaged over $20$ trials and the shaded region is the $95\%$ confidence interval. Lower and less erratic is better.
On the dmc dataset, we find that krope reliably produces stable and low msve during the full course of training.
@@ C.3.2 Learning Graphs for Offline Policy Evaluation on dmc and d4rl Datasets
On the d4rl dataset, we reach the similar conclusions: krope is effective in producing stable and accurate ope estimates. However, in $3/9$ instances, krope does diverge. This divergence is likely related to the discussion in Section 5 and Appendix C.3.1. Recall that krope is a semi-gradient method, which does not optimize any objective function and is susceptible to divergence (Feng et al., 2019; Sutton & Barto, 2018). So while krope representations stabilize value function learning, krope’s learning algorithm may diverge and not converge to krope representations. However, regardless of this result, krope does improve the stability and accuracy of fqe in all cases.
@@ C.3.3 Stability-related Analysis on Custom Datasets
Hyperparameter Sensitivity.
@@ C.3.4 Using fqe Directly for ope
For the sake of completeness, we present results of fqe as an ope algorithm where the action-value network is directly used to estimate the performance of ${\pi_{e}}$. We present the results in Figures 9 and  10. As done in Section 4, we evaluate the performance of fqe and krope based on how they shape the penultimate features of the action-value network. However, when conducting ope, we evaluate two variants: 1) using lspe (-l) and 2) using the same end-to-end fqe action-value network (-e2e).
@@
Regardless of fqe’s hyperparameter sensitivity, it is still interesting to observe that when fqe is used as an ope algorithm, it produces reasonably accurate ope estimates. It even outperforms the fqe+krope combination. The primary difference between using fqe for ope vs. fqe features and lspe for ope is in how the last linear layer is trained. The former is trained by gradient descent while the latter is trained with the iterative lspe algorithm on the fixed features. An interesting future direction will be to explore the learning dynamics of these two approaches.
On a related note, we point out that the training dynamics of fqe are still not well-understood. For example, (Fujimoto et al., 2022) show that the fqe loss function poorly correlates with value error. That is, the fqe loss can be high but value error (and ope error) can be low.